Straight-Line, Declining-Balance, and Sum-of-the-Years'-Digits Methods

A light truck is purchased on January 1 at a cost of $38,730. It is expected to serve for eight years and have a salvage value of $5,690. Calculate the depreciation expense for the first and third years of the truck’s life using the following methods. If required, round your answers to the nearest cent.

Depreciation Expense
Year 1 Year 3
1. Straight-line $ $
2. Double-declining-balance $ $
3. Sum-of-the-years'-digits $ $
Answer #1
Depreciation Expense
Year 1 Year 3
1. Straight-line $4,130 $ 4,130
2. Double-declining-balance $ 9,682.50 $ 5446.41
3. Sum-of-the-years'-digits $7342.22 $ 5506.67

Working

Straight line Method
A Cost $        38,730.00
B Residual Value $           5,690.00
C=A - B Depreciable base $        33,040.00
D Life [in years left ]                             8
E=C/D Annual SLM depreciation $           4,130.00
Depreciation schedule-Straight line method
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
1 $          38,730.00 $           4,130.00 $         4,130.00 $      34,600.00
2 $          34,600.00 $           4,130.00 $         8,260.00 $      30,470.00
3 $          30,470.00 $           4,130.00 $      12,390.00 $      26,340.00
Double declining Method
A Cost $        38,730.00
B Residual Value $           5,690.00
C=A - B Depreciable base $        33,040.00
D Life [in years] 8
E=C/D Annual SLM depreciation $           4,130.00
F=E/C SLM Rate 12.50%
G=F x 2 DDB Rate 25.00%
Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $          38,730.00 25.00% $         9,682.50 $         9,682.50 $      29,047.50
2 $          29,047.50 25.00% $         7,261.88 $      16,944.38 $      21,785.63
3 $          21,785.63 25.00% $         5,446.41 $      22,390.78 $      16,339.22
Sum-of-the-Years'-Digits Method
A Cost $        38,730.00
B Residual Value $           5,690.00
C=A - B Depreciable base $        33,040.00
Sum of digits with 5 years (8+7+6+5+4+3+2+1) 36
Year Depreciable base Formula Depreciation expense
1 $          33,040.00 (([33040x8/36]/) $         7,342.22
2 $          33,040.00 (([33040x7/36]/) $         6,424.44
3 $          33,040.00 (([33040x6/36]/) $         5,506.67